# Phoenix: Pauli-Based High-Level Optimization Engine for Instruction Execution on NISQ Devices

### v4

@@ Abstract
Variational quantum algorithms (VQA) based on Hamiltonian simulation represent a specialized class of quantum programs well-suited for near-term quantum computing applications due to its modest resource requirements in terms of qubits and circuit depth. Unlike the conventional single-qubit (1Q) and two-qubit (2Q) gate sequence representation, Hamiltonian simulation programs are essentially composed of disciplined subroutines known as Pauli exponentiations (Pauli strings with coefficients) that are variably arranged. To capitalize on these distinct program features, this study introduces Phoenix, a highly effective compilation framework that primarily operates at the high-level Pauli-based intermediate representation (IR) for generic Hamiltonian simulation programs. Phoenix exploits global program optimization opportunities to the greatest extent, compared to existing SOTA methods despite some of them also utilizing similar IRs. Experimental results demonstrate that Phoenix outperforms SOTA VQA compilers across diverse program categories, backend ISAs, and hardware topologies.
@@ I Introduction
Conventionally, a Pauli exponentiation is synthesized into a 1Q rotation $Z(\theta)$ sandwiched by a pair of symmetric $\mathrm{CNOT}$ trees, subsequently conjugated by some $H$ and $S$ gates. Existing state-of-the-art (SOTA) compilers primarily exploit gate cancellation opportunities exposed by the already synthesized subcircuits, whether by means of the abstract ZX diagram [8, 9, 10] representation or the variants of $\mathrm{CNOT}$ trees [11, 12]. Despite utilizing the Pauli-based IR, the optimization process they formulate is limited to subcircuits and local IR patterns. Furthermore, it assumes the conventional $\mathrm{CNOT}$-based quantum instruction set architecture (ISA). We instead present that a set of Pauli exponentiations can be simultaneously simplified through appropriate Clifford transformations. For example, a list of weight-3 Pauli strings $[ZYY;\,ZZY;\,XYY;\,XZY]$ can be simplified into a weight-2 Pauli string list through the conjugation of a 2Q Clifford operator $C=(H\otimes S)\,\mathrm{CNOT}\,(H\otimes S^{\dagger})$. This approach unlocks greater optimization opportunities entirely at the level of Pauli-based IR while remaining agnostic to the quantum ISA being employed.
@@
pipeline to compile real-world VQA programs into basic quantum gates. It is ISA-independent, routing-aware, and tailored to generic VQA programs such as molecular simulation involving heterogeneous-weight Pauli strings and QAOA including only weight-2 Pauli strings. Differing from existing SOTA methods, Phoenix utilizes an alternative formal description for Pauli-based IR and integrates heuristic optimization strategies to achieve global optimization primarily at the high-level semantic layer. We evaluate Phoenix across diverse VQA programs, backend ISAs, and hardware topologies, demonstrating its superior performance over existing SOTA compilers.
@@
II Evaluation
@@ II-B Main results
II-C Hardware-aware compilation
@@
II-D Comparison in diverse ISAs
@@ II-E QAOA benchmarking
II-F Algorithmic error analysis

### v5

@@ Abstract
Variational quantum algorithms (VQA) based on Hamiltonian simulation represent a specialized class of quantum programs well-suited for near-term quantum computing applications due to its modest resource requirements in terms of qubits and circuit depth. Unlike the conventional single-qubit (1Q) and two-qubit (2Q) gate sequence representation, Hamiltonian simulation programs are essentially composed of disciplined subroutines known as Pauli exponentiations (Pauli strings with coefficients) that are variably arranged. To capitalize on these distinct program features, this study introduces Phoenix, a highly effective compilation framework that primarily operates at the high-level Pauli-based intermediate representation (IR) for generic Hamiltonian simulation programs. Phoenix exploits global program optimization opportunities to the greatest extent, compared to existing SOTA methods despite some of them also utilizing similar IRs. Phoenix employs the binary symplectic form (BSF) to formally describe Pauli strings and reformulates IR synthesis as reducing the column weights of BSF by appropriate Clifford transformations. It comes with a heuristic BSF simplification algorithm that searches for the most appropriate 2Q Clifford operators in sequence to maximally simplify the BSF at each step, until the BSF can be directly synthesized by basic 1Q and 2Q gates. Phoenix further performs a global ordering strategy in a Tetris-like fashion for these simplified IR groups, carefully balancing optimization opportunities for gate cancellation, minimizing circuit depth, and managing qubit routing overhead. Experimental results demonstrate that Phoenix outperforms SOTA VQA compilers across diverse program categories, backend ISAs, and hardware topologies.
@@ I Introduction
Figure 1: Conventional Pauli exponentiation synthesis v.s. Simultaneous simplification by Clifford conjugation. (a) Pauli exponentiations are synthesized by gate set $\{H,S,S^{\dagger},Z(\theta),\mathrm{CNOT}\}$, through variable $\mathrm{CNOT}$-tree unrolling schemes. (b) Multiple 3Q Pauli exponentiations can be simultaneously simplified into 2Q Pauli exponentiations, through a 2Q Clifford conjugation, where $C=(H\otimes S)\,\mathrm{CNOT}\,(H\otimes S^{\dagger})$ in this example. (c) The simplified Pauli exponentiations can be rebased to versatile quantum gate sets, such as $\mathrm{CNOT}$-based and $\mathrm{B}$-based ISAs, with significantly reduced 2Q gate count.
Conventionally, a Pauli exponentiation is synthesized into a 1Q rotation $Z(\theta)$ sandwiched by a pair of symmetric $\mathrm{CNOT}$ trees, subsequently conjugated by some $H$ and $S$ gates, as exemplified in Fig. 1 (a). Existing state-of-the-art (SOTA) compilers primarily exploit gate cancellation opportunities exposed by the already synthesized subcircuits, whether by means of the abstract ZX diagram [8, 9, 10] representation or the variants of $\mathrm{CNOT}$ trees [11, 12]. Despite utilizing the Pauli-based IR, the optimization process they formulate is limited to subcircuits and local IR patterns. Furthermore, it assumes the conventional $\mathrm{CNOT}$-based quantum instruction set architecture (ISA). We instead present that a set of Pauli exponentiations can be simultaneously simplified through appropriate Clifford transformations. For example, Fig. 1 (b) shows that the list of weight-3 Pauli strings $[ZYY;\,ZZY;\,XYY;\,XZY]$ can be simplified into a weight-2 Pauli string list through the conjugation of a 2Q Clifford operator $C=(H\otimes S)\,\mathrm{CNOT}\,(H\otimes S^{\dagger})$. This approach unlocks greater optimization opportunities entirely at the level of Pauli-based IR while remaining agnostic to the quantum ISA being employed.
@@
pipeline to compile real-world VQA programs into basic quantum gates. It is ISA-independent, routing-aware, and tailored to generic VQA programs such as molecular simulation involving heterogeneous-weight Pauli strings and QAOA including only weight-2 Pauli strings. Differing from existing SOTA methods, Phoenix utilizes an alternative formal description for Pauli-based IR and integrates heuristic optimization strategies to achieve global optimization primarily at the high-level semantic layer. We evaluate Phoenix across diverse VQA programs, backend ISAs, and hardware topologies, demonstrating its superior performance over existing SOTA compilers. Our main contributions are as follows:
  1. 1.
    We utilize the binary symplectic form (BSF) as the formal description of Pauli-based IR and reformulate the IR synthesis process as simultaneous simplification on BSF by appropriate Clifford conjugations that simultaneously reduce the BSF’s column weights, enabling global optimization to the largest extent while operating in the high-level IR.
  2. 2.
    We propose a heuristic BSF simplification algorithm as the core optimization pass in Phoenix. It iteratively searches for the most appropriate 2Q Clifford operators to quickly lower the weight of the BSF, until the BSF can be directly synthesized by basic 1Q and 2Q gates.
  3. 3.
    We quantify the circuit depth overhead induced by assembling (ordering) simplified IR groups. A uniform cost function is devised to comprehensively account for circuit depth overhead, gate cancellation, and qubit routing overhead. Under the guidance of that, Phoenix integrates a Tetris-like IR group ordering procedure to reliably achieve a good ordering scheme.
@@
II Related Works
Based on ZX diagram representations of Pauli gadgets. Compilers such as TKet [13, 8], PauliOpt [9], and PCOAST [10] resynthesize quantum circuits into its ZX diagram representation. ZX diagrams are used to visually represent and simplify the commutation relations and phase interactions between Pauli operators, facilitating the reduction of circuit depth and gate count through algebraic and graphical transformations. Although it can efficiently resynthesize Pauli gadgets, the commutation rules it leverages occur locally, and it is hard to operate in a hardware-aware manner.
Based on synthesis variants of Pauli-based IR. Dedicated compilers like Paulihedral[11] and Tetris [12] leverage Pauli-based IR to identify gate cancellation opportunities between nearest-neighbor IRs, exposed by the variants of their synthesis schemes based on $\mathrm{CNOT}$-tree unrolling. They proposed sophisticated co-optimization techniques to minimize $\mathrm{CNOT}$ gate for both logical-level synthesis and $\mathrm{SWAP}$-based routing achieve good performance especially on limited-topology NISQ devices. However, their optimization scope is confined to a finite set of local subcircuit patterns, and they rely solely on the $\mathrm{CNOT}$-based ISA.
III BSF and Clifford Formalism
Figure 2: Examples of Clifford transformations on BSF. Columns $a$ and $b$, corresponding to qubit $a$ and $b$, for $X$ and $Z$ blocks are indicated by $x_{a}$, $x_{b}$, $z_{a}$, $z_{b}$, respectively. (a) $H$ gate acting on qubit $a$ will exchanges $x_{a}$ and $z_{a}$. (b) $S$ gate acting on qubit $a$ results in $z_{a}\leftarrow z_{a}\oplus x_{a}$. (c) $C(Z,X)_{a,b}$, i.e., the $\mathrm{CNOT}$ gate, results in $x_{b}\leftarrow x_{b}\oplus x_{a}$ and $z_{a}\leftarrow z_{a}\oplus z_{b}$. (d) $C(X,X)$ transformation is equivalent with successively applying $H$ on $a$, $C(Z,X)_{a,b}$, and $H$ on $a$, resulting in $x_{a}\leftarrow x_{a}\oplus z_{b}$ and $x_{b}\leftarrow x_{b}\oplus z_{a}$.
We establish our synthesis scheme by representing Pauli strings in the binary symplectic form (BSF). In the BSF, each $n$-qubit Pauli string is represented as a row in a tableau, with columns divided into two sections: $[X\,|\,Z]$. Here, $[X_{i,j}\,|\,Z_{i,j}]$ denotes the $j$-th component of the $i$-th Pauli operator. The encoding maps $X$ to [1| 0]delimited-[]conditional1 0[1\,|\,0][ 1 | 0 ], $Z$ to [0| 1]delimited-[]conditional01[0\,|\,1][ 0 | 1 ], $Y$ to [1| 1]delimited-[]conditional11[1\,|\,1][ 1 | 1 ], and $I$ to [0| 0]delimited-[]conditional0 0[0\,|\,0][ 0 | 0 ]. For instance, the Pauli string simplification shown in Fig. 1 (b) follows the formulation
$\left[\begin{array}[]{ccc|ccc}0&1&1&1&1&1\\ 0&0&1&1&1&1\\ 1&1&1&0&1&1\\ 1&0&1&0&1&1\\ \end{array}\right]\xrightarrow{C(X,Y)_{1,2}}\left[\begin{array}[]{ccc|ccc}0&1&0&1&1&0\\ 0&0&0&1&1&0\\ 1&1&0&0&1&0\\ 1&0&0&0&1&0\\ \end{array}\right],$
where $C(X,Y)_{1,2}$ is the Clifford gate in Fig. 1 (b) while written in the universal controlled gate representation
$C(\sigma_{0},\sigma_{1})=\frac{1}{2}((I+\sigma_{0})\otimes I+(I-\sigma_{0})\otimes\sigma_{1}),\,\sigma_{j}\in\left\{X,Y,Z\right\}.$
Any universal controlled gate is equivalent to $\mathrm{CNOT}$ up to local $H$ and $S$ conjugations. For example, $C(Z,X)=\mathrm{CNOT}$, and $C(X,Y)=(H\otimes S)\,\mathrm{CNOT}\,(H\otimes S^{\dagger})$.
By definition, any operator $C$ from the Clifford group is unitary and has the property that $CPC^{\dagger}$ is also a Pauli operator for any Pauli operator $P$. The update rules of some Clifford operators in terms of the BSF are demonstrated in Fig. 2, in which only 2Q Clifford operators (Clifford2Q) potentially have nontrivial effects of reducing the Pauli strings’ weights. For any $C(\sigma_{0},\sigma_{1})$, its tableau update rule is a combinatorial sequence of $H$, $S$ and $C(Z,X)$ (or $C(Z,Z)$) transformations. For example, $C(X,Y)_{a,b}$ exhibits the rule
$[x_{a},\,x_{b}\,|\,z_{a},\,z_{b}]\rightarrow[x_{a}\oplus x_{b}\oplus z_{b},\,z_{a}\oplus z_{b}\,|\,z_{a},\,z_{a}\oplus z_{b}].$ (3)
Therefore, the basic approach of our synthesis scheme is to find an appropriate sequence of Clifford2Q operators to simplify the BSF [^1] The BSF that needs to simplify does not contain any weight-1 Pauli string. until it can be directly synthesized by basic 1Q and 2Q gates, i.e., the quantity “total weight”
[…]
is no more than 2, where $r_{x/z}^{(i)}$ is the $i$-th row of $X$ or $Z$ block of the BSF, and the norm is the sum of all binary entries. The key lies in how to find the Clifford2Q that can reduce the weights of as many Pauli strings as possible within the BSF.
IV Our Proposal: Phoenix
IV-A Overview
In the proposed framework Phoenix, Pauli-based IRs (exponentiations) are first grouped according to the same set of qubit indices non-trivially acted on. Then, the BSF simplification algorithm is applied to the BSF of each IR group, generating the simplified subcircuit composed of $\mathrm{CNOT}$-equivalent Clifford2Q operators and Pauli exponentiations with weights no more than 2. Phoenix further selectively assemble (order) these unarranged simplified IR groups in a Tetris-like style that minimizes a uniform subcircuit assembling cost function that incorporates gate cancellation, circuit depth increase, and qubit routing overhead. By defining appropriate metric functions and heuristic algorithms, Phoenix achieves superior optimization (less 2Q gate count and circuit depth) compared to SOTA methods, efficiently handling compilation for large-scale VQA programs.
IV-B BSF simplification for each IR group
1
Input : Pauli strings list pls
Output : Reconfigured circuit components list cfg
2
[…]
4 while bsf.totalWeight() $>$ 2 do
5       local_bsf $\leftarrow$ bsf.popLocalPaulis();
6
      $C\leftarrow\emptyset$ ;
        // Clifford2Q candidates
       $B\leftarrow\emptyset$ ;
        // Each element of $B$ results from applying each Clifford2Q candidate on bsf
[…]
        // Cost functions calculated on each element of $B$
7       for cg in CLIFFORD_2Q_SET do
8             for i, j in $\textsc{combinations}(\textsc{range}(n),2)$ do
                   cliff $\leftarrow$ cg.on$(i,j)$ ;
                    // qubits acted on
9                   bsf $\leftarrow$ bsf.applyClifford2Q(cliff);
10                   cost $\leftarrow$ calculateBSFCost(bsf);
11                   $C.\textsc{append}$(cliff);
12                   $B.\textsc{append}$(bsf);
13                   costs.append(cost);
14
15             end for
16
17       end for
18      bsf $\leftarrow\textsc{BSFWithMinCost}(B,\textit{costs})$;
[…]
20       cliffs_with_locals.append((cliff, local_bsf));
21
22 end while
23cfg.append(bsf);
24 for cliff, local_bsf in cliffs_with_locals do
       // Clifford2Q operators are added as conjugations, with local Pauli strings peeled before each epoch
25       cfg.prepend(cliff);
26       cfg.append(local_bsf);
27       cfg.append(cliff);
28
29 end for
30
Algorithm 1 Pauli Strings Simplification in BSF
To search for appropriate Clifford2Q operators in BSF simplification, it suffices to just focus on a set of Clifford2Q generators of the 2Q Clifford group. We choose the six universal controlled gates
$\left\{C(X,X),C(Y,Y),C(Z,Z),C(X,Y),C(Y,Z),C(Z,X)\right\}$ (5)
that are independent of each other and span the entire 2Q Clifford group [15]. It is a natural choice of sets of Clifford2Q generators, as each of them is Hermitian and equivalent to $\mathrm{CNOT}$, and its corresponding BSF’s tableau update rule can be defined as a combination of the update rules for $H$, $S$, and $\mathrm{CNOT}$.
To determine which Clifford2Q operator to select from Eq. 5, we define a heuristic cost function
$\coloneqqsubscriptcostbsf\coloneqq\displaystyle\mathrm{cost}_{\mathrm{bsf}}\coloneqq\,$ $w_{\mathrm{tot.}}*n_{\mathrm{n.l.}}^{2}+\sum\nolimits_{\langle i,j\rangle}\lVert r_{x}^{(i)}\lor r_{z}^{(i)}\lor r_{x}^{(j)}\lor r_{z}^{(j)}\rVert$
$\quad+\frac{1}{2}\sum\nolimits_{\langle i,j\rangle}(\lVert r_{x}^{(i)}\lor r_{x}^{(j)}\rVert+\lVert r_{z}^{(i)}\lor r_{z}^{(j)}\rVert)$ (6)
to quantify the disparity of the current BSF from a desired BSF that requires no further simplification ($w_{\mathrm{tot.}}\leq 2$). $\mathrm{cost}_{\mathrm{bsf}}$ is the combined weight overlap of both $X$-part and $Z$-part among each pair of Pauli strings of a BSF, with a bias considering the impact of the number of nonlocal (n.l., that is, weight larger than 1) Pauli strings.
Algorithm 1 illustrates the BSF simplification procedure in detail. It takes set of Pauli strings as input and outputs a configuration sequence cfg, in which each component is either a Clifford2Q generator from Eq. 5 or a BSF with $w_{\mathrm{tot.}}$ at most 2. Specifically, at each Clifford2Q search epoch, Crefalgo:simplification identifies the Clifford 2Q generator from Eq. 5 and qubit pair to act upon that leads to the greatest reduction in the cost function defined in Eq. 6, and the BSF is updated accordingly. This greedy process continues iteratively until the BSF’s $w_{\mathrm{tot.}}$ is no more than 2. Before each search epoch, local Pauli strings are peeled from the BSF, as they represent 1Q Pauli rotations and do not induce synthesis overhead. The output cfg represents a simplified IR group, while still in high-level semantics (Clifford2Q, 1Q Pauli-based IRs, and successive 2Q Pauli-based IRs), independent of any specific quantum ISA.
In theory, simplifying a BSF is always achievable, for example, we can reduce an individual Pauli string to be weight-1, peel it off, and then iterate. The efficacy of Algorithm 1 lies in its simultaneous simplification mechanism guided by the cost function in Eq. 6.
IV-C Ordering IR groups in a Tetris-like style
Figure 3: Tetris-like circuits assembling. (a) $e_{l}$ and $e_{r}$ examples. (b) Scenario I: $\mathrm{cost}_{\mathrm{depth}}=\textsc{sum}(e_{r}+e_{l}^{\prime})$; Scenario II: $\mathrm{cost}_{\mathrm{depth}}=\textsc{sum}(e_{r}+e_{l}^{\prime}-1)$.
With simplified IR groups compiled via BSF simplification, Phoenix further performs a Tetris-like IR group ordering procedure that aims to minimize a uniform assembling cost metric. This is unlike [11] and [12] where the ordering is primarily determined by gate cancellation opportunities between IR groups. We can abstract the circuit structure exhibited by each simplified IR group into something resembling a Tetris block. Then, we choose the ordering based on the circuit-depth cost function of assembling Tetris blocks, with gate cancellation opportunities and qubit routing overhead also taken into account. We primarily have three ingredients:
  1. 1.
    Endian vectors of circuits and the depth cost function. As shown in Fig. 3 (a), we define a pair of vectors $e_{l}$ and $e_{r}$ for each subcircuit (which would correspond to a simplified IR group in our case). The $i$-th entry of $e_{l}$ ($e_{r}$) refers to how many layers one has to traverse starting from the left (right) side before qubit $i$ is acted upon. The layers are defined by grouping neighboring 2Q gates that act on different qubits. See Fig. 3 (a) for an example of a generic (not necessarily simplified IR group) circuit. With this definition, the depth overhead for assembling two subcircuits—the proceeding subcircuit with left-endian vectors $e_{r}$ and the succeeding subcircuit with left-endian vector $e_{l}^{\prime}$ is given by
[…]
    See Fig. 3 (b) for an example.
  2. 2.
    Clifford2Q cancellation. The simplified IR group usually exposes Hermitian Clifford2Q operators at the two ends of the subcircuit. Therefore some Clifford2Q gate cancellation opportunities could be exploited, which may or may not decrease overall circuit depth, as depicted in Fig. 4 (a). We can add this consideration into $\mathrm{cost}_{\mathrm{depth}}$: if $m$ pairs Clifford2Q cancelled, while (a) without decreasing subcircuit depth, $\mathrm{cost}_{\mathrm{depth}}\leftarrow\mathrm{cost}_{\mathrm{depth}}-2m$; (b) with one side’s subcircuit depth decreases, $\mathrm{cost}_{\mathrm{depth}}\leftarrow\mathrm{cost}_{\mathrm{depth}}-2m-n$; (c) with both sides’ subcircuit depths decrease, $\mathrm{cost}_{\mathrm{depth}}\leftarrow\mathrm{cost}_{\mathrm{depth}}-2m-2n$.
  3. 3.
    Consideration of qubit routing overhead. We try to mitigate the routing overhead for hardware-aware compilation by comparing the qubit interaction graphs of subcircuits. This approach is not limited to specific routing schemes (e.g. three-$\mathrm{CNOT}$ unrolling of $\mathrm{SWAP}$-based routing, ancilla-based bridge gate [16]) and hardware topologies. Intuitively, two subcircuits with more “similar” qubit interaction behaviors require less mapping transition overhead between them, as demonstrated by Fig. 4 (b). Therefore, we introduce a factor characterized by the similarity between the qubit interaction graphs of two subcircuits in the cost function:
    $\mathrm{cost}_{\mathrm{depth}}\leftarrow\frac{1}{s}\cdot\mathrm{cost}_{\mathrm{depth}},\quad s=\sum\nolimits_{i}\frac{\langle D_{i},D_{i}^{\prime}\rangle}{\lVert D_{i}\rVert_{2}\lVert D_{i}^{\prime}\rVert_{2}},$ (7)
    where $D$ ($D^{\prime}$) is the distance matrix of the preceding (succeeding) subcircuit’s qubit interaction graph of the tail (head) part. The tail (head) is defined as starting from the right (left) of the subcircuit and incorporating more and more 2Q gates until all qubits are acted upon. $D_{i,j}$ represents the shortest path length between qubit $i$ and qubit $j$. $D_{i}$ is the $i$-th row of $D$.
Figure 4: Gate cancellation opportunities and routing-aware assembling. (a) Clifford2Q cancellation between the preceding and succeeding subcircuits may or may not induce circuit depth decrease. (b) The subcircuit (upper right) whose qubit interaction graph is more similar to that of the already assembled subcircuit (left) is preferred over the other (lower right).
In practice, these IR groups are first pre-arranged in descending order of their weight (subcircuit width). Phoenix looks ahead for a certain number of subcircuits to find the one with the minimum $\mathrm{cost}_{\mathrm{depth}}$ with respect to the last assembled subcircuit. This process iterates until all subcircuits are assembled.
V Evaluation
@@ V-B Main results
V-C Hardware-aware compilation
@@
V-D Comparison in diverse ISAs
@@ V-E QAOA benchmarking
V-F Algorithmic error analysis
@@ VI Conclusion and Outlook
Acknowledgement
This work is supported by AI Chip Center for Emerging Smart Systems (ACCESS), sponsored by InnoHK funding, Research Grants Council of HKSAR (No. 16213824) and National Key Research and Development Program of China (No. 2023YFA1009403), National Natural Science Foundation of China (No. 12347104), Beijing Natural Science Foundation (No. Z220002). DD would like to thank God for all of His provisions.